OpenStudy (is3535):

A sack holds 6 red delicious apples and 9 honey crisp apples. An apple is pulled out of the sack at random and then placed back in the sack. How many times can you expect that a red delicious apple will be pulled out of the sack in 80 tries? A. 28 B. 32 C. 40 D. 53

OpenStudy (is3535):

c - 40. You use probability. 6+9 = 15 There are 6 red delicious apples in the sack. Since you want to find the probability of pulling a red delicious apple from the sack, you would divide. 6 \frac{6}{15} = .40 .40 out of 1 would be 40 apples. Since the variables are independent, the probability will be the same each time you pull an apple from the sack.

OpenStudy (anonymous):

here are 15 total right? and 6 red compute 615×80 or even 25×80 or even .4×80 any way you like

OpenStudy (is3535):

40

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

Each time you draw, you can consider that you get 0.4 of 1 apple. Therefore if you make 80 draws you get 0.4 * 80 = 32 apples.